# Barron Space Representations for Elliptic PDEs with Homogeneous Boundary Conditions

### v1

@@ 3.2 Step 2: Barron Norm Estimates
[…]
@@
    Under Assumptions 2.1 and 2.11, the iterative scheme (3.8) for the Dirichlet problem (1.3) satisfies
@@
    Under Assumptions 2.1 and 2.12, the iterative scheme (3.9) for the Neumann problem (1.4) satisfies
@@ 3.2.1 Sobolev Gradient Flow via the Inverse Elliptic Operator
$(I-\Delta)w=\mathcal{L}u-f\quad\text{in }\Omega,\quad w=0\quad\text{on }\partial\Omega.$
This implies that the energy gradient for the Dirichlet problem can be expressed as
@@ 3.2.2 Analysis of the Barron Norm
(3). We prove only the first inequality, as the second is analogous. Let
@@
A similar estimation as in (2.3) shows that both series $\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$ and $\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}C_{k^{\prime}}(x)$ converge absolutely and uniformly on $\Omega$. Hence, by Fubini’s theorem, the product $gh$ can be computed via termwise multiplication:
$gh$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\prod_{j=1}^{d}\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
@@
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\pm\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{v_{m}},$
where each $v_{m}=(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime})$ denotes one of the $4^{d}$ possible sign combinations such that $v_{m}\in\mathbb{N}^{d}$, which is ensured by the choice of signs in the last equality.
Using the fact that $\|v_{m}\|=\|\pm k\pm k^{\prime}\|\leq\|k\|+\|k^{\prime}\|$, we estimate
$1+\pi^{n}\|v_{m}\|^{n}\leq 1+\pi^{n}\left(\|k\|+\|k^{\prime}\|\right)^{n}\leq\left(1+\pi^{n}\|k\|^{n}\right)\left(1+\pi^{n}\|k^{\prime}\|^{n}\right).$ (3.24)
By the reordering invariance of the Barron norm stated in Remark 2.9 and estimate (3.24), we obtain
$\|gh\|_{\mathcal{B}_{s}^{n}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\pm\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{v_{m}}(x)\right\|_{\mathcal{B}_{s}^{n}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\frac{1}{2^{d}}|a_{k}b_{k^{\prime}}|\left(1+\pi^{n}\|v_{m}\|^{n}\right)$
@@
(4). We prove only the first identity, as the second follows analogously. Let
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x).$
Combining Proposition 3.10 with the following three estimates, valid for each $i=1,\dots,d$,
$\sum_{k\in\mathbb{N}^{d}}\left|a_{k}\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}<\infty,$
$\sum_{k\in\mathbb{N}^{d}}\left|a_{k}\frac{\pi k_{i}}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\frac{\cos(\pi k_{i}x)}{\sin(\pi k_{i}x)}\right|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}<\infty,$
$\sum_{k\in\mathbb{N}^{d}}\left|a_{k}\frac{-\pi^{2}k_{i}^{2}}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}<\infty,$
we conclude that the following series converge absolutely and uniformly on $\Omega$ by the Weierstrass $M$-test:
$\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x),\quad\partial_{i}\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x),\quad\partial_{ii}\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x).$
Therefore, we may freely exchange the order of summation and differentiation. In particular, we have
$(I-\Delta)_{D}\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x).$
Since the right-hand side belongs to $L^{2}(\Omega)$, the uniqueness of $(I-\Delta)_{D}^{-1}$ implies
$(I-\Delta)_{D}^{-1}\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x),$
and hence,
@@
$\left\|(I-\Delta)_{D}^{-1}g\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}|a_{k}|=\frac{1}{2}\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}.$
@@
We prove the case under Assumption 2.11; the case under Assumption 2.12 follows by the same argument.
@@
Since
$\sum_{k\in\mathbb{N}^{d}}|a_{k}\partial_{i}C_{k}(x)|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}\pi k_{i}|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi\|k\|\right)=\|g\|_{\mathcal{B}_{c}^{1}(\Omega)}<\infty,$
it follows from the Weierstrass $M$-test that the series $\sum_{k\in\mathbb{N}^{d}}a_{k}\partial_{i}C_{k}(x)$ converges absolutely and uniformly on $\Omega$. Therefore, we are justified in interchanging the order of summation and differentiation. A similar argument applies to the series $\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}\,\partial_{i}S_{k^{\prime}}(x)$, which also converges absolutely and uniformly. Moreover, since both series are absolutely convergent, we may apply Fubini’s theorem to freely interchange the order of summation when taking products of their terms. Thus, we compute
$\partial_{i}A_{ii}\partial_{i}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(\partial_{i}C_{k}\partial_{i}S_{k^{\prime}}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)\pm\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ii}\partial_{i}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}(I-\Delta)^{-1}_{D}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{i}^{\prime}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
@@
By a similar argument of absolute and uniform convergence as in the proof of (3.25), the following computation is justified:
$\partial_{i}A_{ij}\partial_{j}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\quad\quad\cdot\cos\left(\pi k_{i}x_{i}\right)\sin\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)\pm\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ij}\partial_{j}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}(I-\Delta)^{-1}_{D}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{j}^{\prime}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
@@
By a similar argument of absolute and uniform convergence as in the proof of (3.25), the following computation is justified:
$A_{ii}\partial_{ii}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(C_{k}\partial_{ii}S_{k^{\prime}}\right)$
@@
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ii}\partial_{ii}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}(I-\Delta)^{-1}_{D}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$

### v2

@@ 3.2 Step 2: Barron Norm Estimates
$p_{D}(d)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,D}d^{2}+\alpha_{\ast}\ell_{c}+1,\quad p_{N}(d)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,N}d^{2}+\alpha_{\ast}\ell_{c}+1.$ (3.19)
@@
    Under Assumptions 2.1 and 2.5, the iterative scheme (3.6) for the Dirichlet problem (1.3) satisfies
@@
    Under Assumptions 2.1 and 2.6, the iterative scheme (3.7) for the Neumann problem (1.4) satisfies
@@ 3.2.1 Sobolev Gradient Flow via the Inverse Elliptic Operator
$(I-\Delta)w=\mathcal{L}u-f\quad\text{in }\Omega,\quad w=0\quad\text{on }\partial\Omega,$
where
$\mathcal{L}u\coloneqq-\nabla\cdot A\nabla u+cu$
is the second-order elliptic operator. This implies that the energy gradient for the Dirichlet problem can be expressed as
@@ 3.2.2 Analysis of the Barron Norm
(3). We prove only the first inequality, as the second one is analogous. By the standard uniform convergence argument in analysis, the interchange of summation when taking products in the following proofs is justified, and the details will be omitted.
Let
@@
We compute
$g(x)h(x)$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\prod_{j=1}^{d}\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
@@
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{k+\epsilon\circ k^{\prime}}(x).$
We estimate
$1+\pi^{n}\|k+\epsilon\circ k^{\prime}\|^{n}\leq 1+\pi^{n}\left(\|k\|+\|k^{\prime}\|\right)^{n}\leq\left(1+\pi^{n}\|k\|^{n}\right)\left(1+\pi^{n}\|k^{\prime}\|^{n}\right).$
Thus, we obtain
$\|gh\|_{\mathcal{B}_{s}^{n}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{n}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{1}{2^{d}}|a_{k}b_{k^{\prime}}|\left(1+\pi^{n}\|k+\epsilon\circ k^{\prime}\|^{n}\right)$
@@
(4). We prove only the first identity, as the second follows analogously. Some standard arguments imply that $(I-\Delta)_{D}^{-1}$ and the summation are interchangeable, i.e.,
@@
$\left\|(I-\Delta)_{D}^{-1}g\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}=\sum_{k\in\mathbb{N}^{d}}|a_{k}|=\frac{1}{2}\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}.$
@@
We prove the case under Assumption 2.5; the case under Assumption 2.6 follows by the same argument. By the standard uniform convergence argument in analysis, the interchange of the order of summation and differentiation, as well as the interchange of summation when taking products in the following proofs, is justified. We will omit the details when applying these arguments.
@@
We compute
$\partial_{i}A_{ii}\partial_{i}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(\partial_{i}C_{k}\partial_{i}S_{k^{\prime}}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\qquad\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\cdot\frac{1}{2^{d}}\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i\end{subarray}}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$\qquad\cdot\left(\sin\left(\pi(k_{i}+k_{i}^{\prime})x_{i}\right)+\sin\left(\pi(k_{i}-k_{i}^{\prime})x_{i}\right)\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\epsilon_{i}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ii}\partial_{i}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\epsilon_{i}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\epsilon_{i}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{i}^{\prime}$
$\leq$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
@@
We compute
$\partial_{i}A_{ij}\partial_{j}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\qquad\cdot\cos\left(\pi k_{i}x_{i}\right)\sin\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq j\end{subarray}}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$\qquad\cdot\left(\sin\left(\pi(k_{j}+k_{j}^{\prime})x_{j}\right)+\sin\left(\pi(k_{j}-k_{j}^{\prime})x_{j}\right)\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ij}\partial_{j}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{j}^{\prime}$
$\leq$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
@@
We compute
$A_{ii}\partial_{ii}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(C_{k}\partial_{ii}S_{k^{\prime}}\right)$
@@
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ii}\partial_{ii}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$